Circular No. 7107

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998ee IN ANONYMOUS GALAXY F. Patat, European Southern Observatory (ESO); M. Salvo, Padova

University; and A. Pizzella, ESO, report on their observations of SN 1998ee (cf. IAUC 7030) with the ESO 1.5-m telescope (+ Boller & Chivens spectrograph; range 320–1000 nm, resolution 1.2 nm) on Feb. 8.19 UT at La Silla: “A full reduction of the CCD spectrogram shows that this is a peculiar type-II supernova. The only detectable features are Hα and Hβ emissions, superimposed on a slightly blue continuum. The expansion velocities, deduced from their FWHM, are 5900 and 5400 km/s, respectively. On top of the Hα emission, a clear unresolved P-Cyg profile is visible, possibly accompanied by other weaker features. The same structure, even though less pronounced, is present also in the Hβ line. This indicates the probable presence of a dense expanding medium around the supernova, similar to what has been seen in SN 1994aj (Benetti et al. 1998, MNRAS 294, 448). The expansion velocity of this material, computed from the minimum of the absorption through, is ∼ 1000 km/s. The recession velocity of the parent galaxy, deduced from the unresolved Hα emission, is 14 900 km/s. Assuming that the supernova was discovered around maximum, when B ∼ 17.5 (IAUC 7030), it reached an absolute B magnitude of −19.3 (H_{0} = 65 km s^{−1} Mpc^{−1}), which is definitely brighter than normal type-II supernovae.”

SUPERNOVA 1999Z IN UGC 5608 S. Jha, P. Garnavich, P. Challis, and R. Kirshner, Harvard-Smithson-

ian Center for Astrophysics, report that a spectrum of SN 1999Z was taken by M. Calkins on Feb. 9.3 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph) and shows it to be a type-IIn supernova. The supernova exhibits narrow (FWHM < 300 km/s) Balmer emission lines with somewhat broader (FWHM 1600 km/s) bases. The recession velocity of the host galaxy is 15 100 km/s, derived using narrow Hα emission from a spiral arm. The supernova emission features are observed at a redshifted velocity of 15 600 km/s. Concurrent images taken by P. Zhao with the Whipple Observatory 1.2-m telescope (+ AndyCam) yield an estimated magnitude V = 17.4.

Corrigendum. On IAUC 7106, line 5, for 3′′.6 west and 9′′ south read 11′′ west and 2′′ south

1999 February 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green